The longest side of the triangular section is
8 feet shorter than twice the shortest side. The third side is
4 feet longer than the shortest side. The perimeter is
60 feet. How long is each​ side?

Answer by addingup(3677)

a = 2b-8
c = b+4
and
a+b+c = 60
substitute with the values above:
2b-8+b+4+b = 60
4b-4 = 60
4b = 64
b = 16
~~~~~~~~~~~~~~~~~~~~~~~~~
So now we have that:
b = 16
a = 2b-8 = 2(16)-8 = 24
c = b+4 = 16+4 = 20
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~
Check:
24+16+20 = 60 Correct